# Estimated Heating Rates Due to Cyclotron Damping of Ion-scale Waves Observed by Parker Solar Probe

### v2

@@ 4.1 $Q_{p}/(Q_{p}+Q_{e})$ consistent with the dominant proton heating that is observed near the Sun
Because cyclotron damping onto electrons is negligible, we evaluate $Q_{p}/(Q_{p}+Q_{e})$ by considering just $Q_{e,Landau}$ for electrons while we consider both $Q_{p,cyc}$ and $Q_{p,Landau}$ for protons. Moreover the estimations of $Q_{p/e,Landau}$ and $Q_{p,cyc}$ are possible in intervals where the cascade model employed in SKM23 is accurate and waves are present, respectively. Table 2 shows the percentages of the analyzed intervals where $Q_{p/e,Landau}$ and $Q_{p,cyc}$ estimations are available.
@@
Waves present 26.54 46.06
absent/no overlap 16.04 11.36
@@
We consider intervals where both $Q_{p/e,Landau}$ and $Q_{p,cyc}$ available (i.e. 26.54% of intervals, green shaded in Table 2) and evaluate moving averages of $Q_{p}$ and $Q_{e}$ and then evaluate $Q_{p}/(Q_{p}+Q_{e})$ (shown in panel (a) of Figure 10) corresponding to $Q_{p,cyc}$ (medium green dash-dotted), $Q_{p,Landau}$ (light green dashed), and their sum (dark green solid) as functions of heliocentric distance. We infer that cyclotron damping greatly enhances the dissipation onto protons, producing much higher rates than those possible by Landau damping alone. By using PSP observations, Bandyopadhyay et al. (2023) have extended to near Sun distances the $Q_{p}/(Q_{p}+Q_{e})$ estimation made by Cranmer et al. (2009) using Ulysses and Helios observations. They estimate that protons receive $\sim$ 80% of plasma heating (cyan in panel (a) of Figure 10). Our estimates of $Q_{p}/(Q_{p}+Q_{e})$ (dark green) are comparable to theirs except in a narrow region between 0.16 and 0.18 au. We hypothesize that this inconsistency occurs because this region has both extremely low $\beta_{p}$ for E1 and E2 observations (see panel (c)), thus inhibiting the proton Landau damping as the protons don’t have large enough thermal velocities for a significant population to undergo Landau resonance with kinetic Alfvén waves, as well as relatively low wave energy (panel (b)), suppressing the cyclotron heating leading to inadequate proton heating via both mechanisms included in our model. These short regions of low-$\beta$ values are likely statistically anomalous, and we expect the plasma parameters to revert to a mean radial distribution (Maruca et al., 2023) with the inclusion of data from additional orbits. We expect this reversion to produce proton heating rates comparable to the elevated levels in nearby distances. This extension will be considered in future work.
[…]
@@ Appendix B Identification of modes in PLUME
    For the three propagating modes, the variation in the parallel group velocity, $v_{g,\parallel}=\frac{\partial\omega}{\partial k_{\parallel}}$ offers a measure to distinguish between solutions. The values of $v_{g,\parallel}$ increase, decrease, and remain roughly constant for the fast, Alfvén, and slow waves, respectively with increasing $k_{\parallel}$; see Equations 8, 14, 16, and 18 in Howes et al. (2014) for analytic approximations of $\omega(k)$ for Alfvén and FMWs at inertial and parallel kinetic scales. At small parallel scales, the ICWs damp heavily onto ions, resulting in a decrease in $v_{g,\parallel}$, while the FMWs don’t damp onto ions; their frequency scales as $\propto kk_{\parallel}$ resulting in an increase in $v_{g,\parallel}$. We calculate $\Delta v_{g,\parallel}$, the change in the parallel group velocity between $k_{\parallel}\rho_{p}=0.003$ and $k_{\parallel}\rho_{p}=1$ for the three modes. The mode with the maximum positive $\Delta v_{g,\parallel}$ value is identified as the FMW, and the one with maximum negative $\Delta\frac{\partial\omega}{\partial k_{\parallel}}$ value is identified as the Alfvén mode.
@@ Appendix C Quantifying circular Polarization of waves
In the plasma frame, magnetic helicity quantifies the rotation of B with respect to space at a fixed time while polarization quantifies the rotation of B with respect to time at a fixed spatial point (Narita et al., 2009). Magnetic helicity is defined with respect to the wavevector, k, while polarization is not. Hence, in the plasma rest frame, magnetic helicity (polarization) has the opposite (same) sign for counter-propagating waves with the same intrinsic polarization. This difference is depicted in Figure 11 which is a variation of Figure 1 from Narita et al. (2009) (where a RH wave is depicted, as opposed to the LH wave in our example). Panels (a1) and (b1) show forward and backward propagating example ICWs in the plasma rest frame where black arrows show the direction of wave propagation. Here XYZ is a right coordinate system where the mean magnetic field, $\mathbf{B_{o}}$ is in the Z direction. The position of $\mathbf{\delta B_{\perp}}=\mathbf{B_{x}}+\mathbf{B_{y}}$ in the plane perpendicular to $\mathbf{B_{o}}$ is shown as colored arrows at 5 consecutive points along Z. When $\mathbf{\delta B_{\perp}}$ points towards X or Y directions, the arrows are colored red or green, respectively.
@@
When the $\mathbf{\delta B_{\perp}}$ of waves is observed with respect to time at a fixed spatial point, the observer sees it in the order t1-t5 for both cases and the rotation of $\mathbf{B_{\perp}}$ (as shown in panels (a4,b4)) is observed to be proton-resonant (LH) for both the waves. The corresponding oscillations of $B_{x}$ and $B_{y}$ with respect to time (shown in panels (a3,b3)) are identical which is consistent with the sign of polarization for both the forward and backward ICWs being the same. On the other hand, panels (a2,b2) show the oscillations of $B_{x}$ and $B_{y}$ with respect to space (z) at a fixed time, and the variation of $B_{y}$ relative to $B_{x}$ is inverted for the forward and backward waves which is consistent with the sign of magnetic helicity being reversed.

### v3

@@ 4.1 $Q_{p}/(Q_{p}+Q_{e})$ consistent with the dominant proton heating that is observed near the Sun
Because cyclotron damping onto electrons is negligible, we evaluate $Q_{p}/(Q_{p}+Q_{e})$ by considering just $Q_{e,Landau}$ for electrons while we consider both $Q_{p,cyc}$ and $Q_{p,Landau}$ for protons. Moreover the estimations of $Q_{p/e,Landau}$ and $Q_{p,cyc}$ are only possible in intervals where the cascade model employed in SKM23 is accurate and waves are present, respectively. Additionally, the cascade model assumes a distribution of turbulent energy with large wavevector anisotropy perpendicular to B. As the angle between $\mathbf{V}_{SW}$ and B decreases, the observed energy spectra is increasingly dominated by any structures parallel to B (e.g. Figure 2 in Horbury et al. (2008)), and insensitive to the perpendicular turbulence. For the comparison of heating rates in this work, we only consider Landau damping estimates in intervals where the perpendicular structure is resolved, $\theta_{BV_{SW}}>30^{\circ}$, discarding approximately half of the intervals where the Landau Damping model was considered accurate in SKM23. Table 2 shows the percentages of the analyzed intervals where $Q_{p/e,Landau}$ and $Q_{p,cyc}$ estimations are available.
@@
Waves present 15.3 55.31
absent/no overlap 12.79 5.88
@@
We consider intervals where both $Q_{p/e,Landau}$ and $Q_{p,cyc}$ available (i.e. 15.3% of intervals, green shaded in Table 2) and evaluate moving averages of $Q_{p}$ and $Q_{e}$ and then evaluate $Q_{p}/(Q_{p}+Q_{e})$ (shown in Figure 10) corresponding to $Q_{p,cyc}$ (medium green dashed-dotted), $Q_{p,Landau}$ (light green dashed), and their sum (dark green solid) as functions of heliocentric distance. We infer that cyclotron damping greatly enhances the dissipation onto protons, producing much higher rates than those possible by Landau damping alone. By using PSP observations, Bandyopadhyay et al. (2023) have extended to near Sun distances the $Q_{p}/(Q_{p}+Q_{e})$ estimation made by Cranmer et al. (2009) using Ulysses and Helios observations. They estimate that protons receive $\sim$ 80% of plasma heating (cyan in panel (a) of Figure 10). Our estimates of $Q_{p}/(Q_{p}+Q_{e})$ (dark green) are in good agreement with theirs.
In 55.31% of intervals, ion-scale waves are observed but the cascade model employed to estimate $Q_{e,Landau}$ is inaccurate. The mean value of $Q_{p,cyc}$ is the same in intervals where the Landau damping model is or is not accurate. In 5.2% of analyzed intervals, there is no overlap between $\omega_{\text{obs}}$ and $\omega_{sc}$ which occurs because of discarding heavily damped PLUME solutions.
@@ Appendix B Identification of modes in PLUME
    For the three propagating modes, the variation in the parallel group velocity, $v_{g,\parallel}=\frac{\partial\omega}{\partial k_{\parallel}}$ offers a measure to distinguish between solutions. The values of $v_{g,\parallel}$ increase, decrease, and remain roughly constant for the fast, Alfvén, and slow waves, respectively with increasing $k_{\parallel}$; see Equations 8, 14, 16, and 18 in Howes et al. (2014) for analytic approximations of $\omega(k)$ for Alfvén and FMWs at inertial and parallel kinetic scales. At small parallel scales, the ICWs damp heavily onto ions, resulting in a decrease in $v_{g,\parallel}$, while the FMWs don’t damp onto ions; their frequency scales as $\propto kk_{\parallel}$ resulting in an increase in $v_{g,\parallel}$. We calculate $\Delta v_{g,\parallel}$, the change in the parallel group velocity between $k_{\parallel}\rho_{p}=0.003$ and $k_{\parallel}\rho_{p}=1$ for the three modes. The mode with the maximum positive $\Delta v_{g,\parallel}$ value is identified as the FMW, and the one with maximum negative $\Delta v_{g,\parallel}$ value is identified as the Alfvén ICW mode.
@@ Appendix C Quantifying circular Polarization of waves
In the plasma frame, magnetic helicity quantifies the rotation of B with respect to space at a fixed time while polarization quantifies the rotation of B with respect to time at a fixed spatial point (Narita et al., 2009). Magnetic helicity is defined with respect to the wavevector, k, while polarization is not. Hence, in the plasma rest frame, magnetic helicity (polarization) has the opposite (same) sign for counter-propagating waves with the same intrinsic polarization. This difference is depicted in Figure 11 which is a variation of Figure 1 from Narita et al. (2009) (where a RH wave is depicted, as opposed to the LH wave in our example). Panels (a1) and (b1) show forward and backward propagating example ICWs in the plasma rest frame where black arrows show the direction of wave propagation. Here XYZ is a right coordinate system where the mean magnetic field, $\mathbf{B_{o}}$ is in the Z direction. The position of $\mathbf{B_{\perp}}=\mathbf{B_{x}}+\mathbf{B_{y}}$ in the plane perpendicular to $\mathbf{B_{o}}$ is shown as colored arrows at 5 consecutive points along Z. When $\mathbf{B_{\perp}}$ points towards X or Y directions, the arrows are colored red or green, respectively.
@@
When the $\mathbf{B_{\perp}}$ of waves is observed with respect to time at a fixed spatial point, the observer sees it in the order t1-t5 for both cases and the rotation of $\mathbf{B_{\perp}}$ (as shown in panels (a4,b4)) is observed to be proton-resonant (LH) for both the waves. The corresponding oscillations of $B_{x}$ and $B_{y}$ with respect to time (shown in panels (a3,b3)) are identical which is consistent with the sign of polarization for both the forward and backward ICWs being the same. On the other hand, panels (a2,b2) show the oscillations of $B_{x}$ and $B_{y}$ with respect to space (z) at a fixed time, and the variation of $B_{y}$ relative to $B_{x}$ is inverted for the forward and backward waves which is consistent with the sign of magnetic helicity being reversed.